OpenStudy (anonymous):

solve: V = π r2h for h ?

OpenStudy (anonymous):

This is the volume of a cylinder, h = height of the cylinder.

OpenStudy (anonymous):

Are you trying to re-write in terms of h?

OpenStudy (anonymous):

yup...agentx5 correct

OpenStudy (compassionate):

h = V/(pi)r^2

OpenStudy (anonymous):

yesss rewrite in terms of h.. ?

mathslover (mathslover):

\[\huge{V=pi * r^2h}\] \[\huge{\frac{V}{pi*r^2}=h}\]

OpenStudy (anonymous):

\[\frac{Volume}{\pi r^2} = h\]

OpenStudy (anonymous):